# SingVisio: Visual Analytics of Diffusion Model for Singing Voice Conversion

### v1

@@ 4.1 Expert Interview
C1: In-Depth Temporal Dynamics Analysis of Diffusion Generation Process
@@
Experts highlighted the importance of visualizing the temporal dynamics of the diffusion generation process in singing voice conversion. The objective is to create detailed visual representations that effectively trace the step-by-step evolution occurring in voice conversion at each diffusion stage. This involves visualizing the progression of various acoustic parameters, such as frequency components and harmonics that evolve over time. The visualizations are expected to provide users with an intuitive understanding of the voice transformation, highlighting the nuanced evolution from a noisy beginning to a structured and coherent output, thereby making the process more transparent and understandable to users without deep technical expertise in machine learning or signal processing.
C2: Comprehensive Performance Metrics Evaluation in Singing Voice Synthesis
@@
This task is centered on tracking evaluation metrics that gauge the quality of the converted voice at each step of the diffusion generation process. These metrics include pitch accuracy, timbre consistency, naturalness, and speech quality. The system should enable a detailed analysis of how each metric evolves with every diffusion step, offering insights into the conversion quality at different phases of the generation process. This comprehensive evaluation is pivotal in identifying aspects where the conversion achieves optimal quality or, conversely, where improvements are needed. This visual representation enhances the interpretability of the evaluation results and fosters insights into the underlying dynamics of the diffusion generation process.
C3: Comparative Analysis of Different Source Singers, Songs, and Target Singers
@@
Through visualization, we aim to systematically compare how different characteristics of source singers, such as vocal tone, pitch range, and singing style, influence the conversion outcome. This will help in identifying specific attributes of source singers that are more amenable to conversion. Additionally, the complexity and structure of the song itself are crucial variables. Songs with intricate melodic lines or complex rhythms might pose more significant challenges in conversion processes. Equally important is the analysis of the target singers’ characteristics. The system should visualize how well the model adapts the source singer’s voice to match the timbre of the target singers. This could lead to valuable insights, such as identifying particularly challenging source-target pairings or songs that consistently yield high-quality conversions. Such analysis is not only crucial for understanding the current model’s performance but also for guiding future improvements and applications in singing voice conversion technology.
These tasks, shaped by expert insights, outline the foundational requirements for our visual analytic system, ensuring that it addresses the key aspects of understanding the diffusion model for singing voice conversion.
@@ 4.2 Analytical Tasks
Our system is a visualization system designed specifically for diffusion-based singing voice conversion tasks. Diffusion-based singing voice conversion itself involves two aspects: in the realm of machine learning, it involves the diffusion generative model, and in the field of audio signal processing, it pertains to singing voice conversion. Therefore, the analytical tasks supported by our visualization system can be divided into two major categories.
In the aspect of machine learning, particularly in the diffusion model, to investigate the evolution and quality of the generated result from each step in the diffusion generation process, our system should support the following two tasks.
@@
Fig. 3: Visual system for diffusion-based singing voice conversion. The system consists of five views. (A) Metric View shows objective evaluation results on the singing voice conversion model, allowing users to interactively explore the performance trend along diffusion steps. (B) Projection View aids users in tracking the data patterns of diffusion steps in the embedding space under different input conditions. (C) Step View provides users with the visualization of Mel spectrogram and pitch contour at one diffusion step. (D) Comparison View facilitates users to compare voice conversion results among different diffusion steps or singers. (E) Control Panel enables users to select various comparison modes and choose different source and target singers to visually understand and analyze the model behavior.
@@ 5 Explainer System
Fig. 3 shows the overview of the explainer system which consists of five components: Control Panel allows users to modify the display mode and choose data for visual comprehension; Step View provides users with an overview of the diffusion generation process; Comparison View makes it easy for users to compare converted results between different conditions; Projection View helps users observe the trajectory of diffusion steps with or without conditions; Metric View displays objective metrics evaluated on the diffusion-based SVC model, enabling users to interactively examine metric trends across diffusion steps.
@@ 5.1 Control Panel
    • 1..1
      Step Comparison This mode primarily focuses on step-wise comparing the diffusion steps in the generation process. It (1) provides an animation of random noise gradually refined for users to have an overview of the whole denoising process in Step View, (2) enables users to flexibly select and detailedly compare the generated results from different diffusion steps in Comparison View.
    • 1..2
      Source Singer Comparison This mode focuses on the pair-wise comparison of converting two different source singers’ audio with the same song to the same target singer. It (1) allows users to select two different source singers, a source song and a target singer, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the two source audio and the target audio in the Comparison View, (3) presents two conversion animations wherein random noise undergoes gradual refinement to transform into the singing voice of the target speaker in the Step View.
    • 1..3
      Song Comparison This mode focuses on the pair-wise comparison of converting two different source audios that are derived from the same singer but contain different songs to the same target singer. It (1) allows users to select a source singer, a target singer but two songs, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the two source singers’ audio and the target singer’s audio in the Comparison View, (3) supplies two conversion animations illustrating the progressive refinement of random noise into the singing voice of the target in the Step View.
    • 1..4
      Target Singer Comparison This mode focuses on the pair-wise comparison of converting the same source singing voice (also means the same song) to two different target singers. It (1) enables users to select a source singer and a source song, but two target singers, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the source singer’s audio, and two target singers’ audio in the Comparison View, (3) provides the two corresponding conversion animations of random noise gradually refined to the target speaker singing voice in the Step View.
    • 1..5
@@
  • 2.
@@
  • 3.
    Projection Embedding A drop-down box to enable users to choose different projection embeddings from different layers. Then, the system displays 2D t-SNE visualization results of the high-dimensional diffusion steps in the Projection View. Specifically, the projection embedding can be the diffusion steps, the combined embeddings of the step and noise, or step, noise and conditions. These embeddings can come from the first, middle, or final residual layer in the diffusion model, as illustrated in Fig. 2.
  • 4.
    Step Controller The Step Controller includes (1) a step slider to smoothly control the diffusion step, (2) a tooltip to display or input a specific step number, and (3) a button named ‘Pin’ enables users to add a specific step’s generated result in the Comparison View.
Fig. 4: Comparison view under the step comparison mode. This view, by default, shows the outcomes of three diffusion steps, including step 999, step 100, and step 10, each marked with a distinct color.
@@ 5.2 Step View
This view enables users to visualize the whole generation process of diffusion in the context of singing voice conversion tasks, which means users can observe how the spectral characteristics change over time as noise is subsequently removed, leading to the desired singing voice conversion. Specifically, it can be observed that the Mel spectrogram transitions from being completely noisy to gradually becoming clearer, and the fundamental frequency curve also transforms from scattered points into a smooth curve. The audio also undergoes a process of gradual optimization from being pure noise to having improved sound quality and intelligibility.
@@ 5.3 Comparison View
To facilitate a more convenient and detailed observation of the intermediate results generated by the diffusion model, we introduce a Comparison View, as shown in Fig. 4. This view provides basic information, including the audible audio, its corresponding Mel spectrogram, and fundamental frequency contour. All the information related to a clip of audio forms a basic block, which we refer to as the ‘basic display unit’. In Fig. 4, there are three basic display units.
@@
On this basic display unit, we can observe the range of the F0 and the pattern of the F0 contour. Through the brush operation, we can synchronously magnify all Mel spectrograms illustrated in this view, thus enabling a more detailed comparison and examination of the spectral differences. When there are multiple basic display units, users can select the checkboxes in the top left corner of any two basic display units. The page will then pop up the visualization of the difference in the Mel spectrogram between these two basic display units, as shown in Fig. 5, allowing for a clearer and more convenient comparison. Specifically, the differences are represented by colors. Warmer colors like reds and oranges signify larger differences, while cooler colors like blues and greens represent smaller differences between the two selected Mel spectrograms.
Fig. 5: The Mel spectrogram difference between two selected Mel spectrograms. When any two Mel spectrograms displayed in the comparison view are selected (via the checkboxes in the top left corner of the basic display unit) (a), a view showing the Mel spectrogram difference will pop up (b).
@@
Furthermore, the components displayed in the Comparison View differs between the Step Comparison Mode and the Condition Comparison Mode. The Step Comparison Mode is primarily used to compare the results of different diffusion steps. In this mode, the Comparison View will display basic display units from three different steps, based on the steps selected by the user. The relative position of different basic display units (corresponding to different steps) can be directly adjusted by dragging. On the other hand, the Condition Comparison Mode sports a similar layout, but mainly compares the results of SVC under different conditions. In this mode, the Comparison View primarily displays the basic display units corresponding to different audios of the source and target singers selected by users. Additionally, in Metric Comparison Mode, the Comparison View illustrates the metric curve over the diffusion step, which is described in Section 5.5.
@@ 5.4 Projection View
High-dimensional diffusion steps can be challenging to interpret directly. t-SNE reduces the dimensionality, projecting the steps into a lower-dimensional space. It allows researchers to gain insights into the intricate structure and relationships within the high-dimensional diffusion step space. By projecting these steps into a lower-dimensional space, t-SNE reveals patterns and trajectories of the diffusion steps, enabling a visual exploration of similarities, differences, and the dynamic evolution of the diffusion process of different generation samples.
Consequently, we design Projection View to present the two-dimensional space obtained by projecting high-dimensional diffusion step embeddings (i.e., the step features in Fig. 2), as shown in Fig. 3. Each point represents a diffusion step, and all 1000 diffusion steps together form a trajectory in space. Users can hover their mouse over the points and slide to inspect the step trajectory. While sliding the mouse, users can simultaneously observe the SVC results transition from a coarse state to a fine state in Step View.
Fig. 6: Projection View zoomed in with ‘Show Step Number’ checked
Additionally, users can enable the “Show Step Number” checkbox to display the step number associated with each step. By scrolling the mouse wheel, they can zoom in or out on the points in the space to explore the distribution of diffusion steps. By clicking on a specific step point, a basic display unit corresponding to the step will be added into the Comparison View. For example, Fig. 6 shows points for three diffusion steps: 100 (blue), 80 (magenta), and 10 (purple).
As described in Section 5.1, the drop-down menu of projection embedding in the control panel provides multiple projection embedding sources, including not only the vanilla diffusion step but also the combination of the diffusion step with noise and condition, as indicated by the red solid dots in Fig. 2. By examining the projection embedding results of combining diffusion step with noise and condition, users can compare the differences in diffusion step trajectories under different condition scenarios.
@@ 5.5 Metric View
The Metric View represents an average of all samples within the testing data pool, providing a comprehensive overview of the model performance with five objective metrics. Upon hovering over a particular metric, the system automatically identifies and selects the best-performing sample. This selection triggers a detailed visualization of the diffusion step for that sample within the Projection View. Additionally, for the chosen sample, the system dynamically computes and displays evaluation metrics, which are then used to plot the “Metric Curve over Diffusion Step” in Comparision View, as shown in Fig. 7.
Fig. 7: The left part is the Metric View with FAD metric selected. The right part is the corresponding “Metric Curve over Diffusion Step” for the best-performing sample on the FAD metric.

### v2

@@ 4.1 Requirement analysis
C1: In-Depth Temporal Dynamics Analysis of Diffusion Generation Process. Experts highlighted the importance of visualizing the temporal dynamics of the diffusion generation process in singing voice conversion. The objective is to create detailed visual representations that effectively trace the step-by-step evolution occurring in voice conversion at each diffusion stage. This involves visualizing the progression of various acoustic parameters, such as frequency components and harmonics that evolve over time. The visualizations are expected to provide users with an intuitive understanding of the voice transformation, highlighting the nuanced evolution from a noisy beginning to a structured and coherent output, thereby making the process more transparent and understandable to users without deep technical expertise in machine learning or signal processing.
@@
C2: Comprehensive Performance Metrics Evaluation in Singing Voice Synthesis. This task is centered on tracking evaluation metrics that gauge the quality of the converted voice at each step of the diffusion generation process. These metrics include pitch accuracy, timbre consistency, naturalness, and speech quality. The system should enable a detailed analysis of how each metric evolves with every diffusion step, offering insights into the conversion quality at different phases of the generation process. This comprehensive evaluation is pivotal in identifying aspects where the conversion achieves optimal quality or, conversely, where improvements are needed. This visual representation enhances the interpretability of the evaluation results and fosters insights into the underlying dynamics of the diffusion generation process.
@@
C3: Comparative Analysis of Different Source Singers, Songs, and Target Singers. Through visualization, we aim to systematically compare how different characteristics of source singers, such as vocal tone, pitch range, and singing style, influence the conversion outcome. This will help in identifying specific attributes of source singers that are more amenable to conversion. Additionally, the complexity and structure of the song itself are crucial variables. Songs with intricate melodic lines or complex rhythms might pose more significant challenges in conversion processes. Equally important is the analysis of the target singers’ characteristics. The system should visualize how well the model adapts the source singer’s voice to match the timbre of the target singers. This could lead to valuable insights, such as identifying particularly challenging source-target pairings or songs that consistently yield high-quality conversions. Such analysis is not only crucial for understanding the current model’s performance but also for guiding future improvements and applications in singing voice conversion technology.
C1 and C2 tasks are related to fundamental knowledge of the diffusion model, which is crucial and beneficial for beginners to understand diffusion models. In contrast, C3 task focuses on exploring the impacts of different conditions on SVC, which is more suitable for experts or researchers seeking an in-depth understanding and analysis of diffusion-based SVC. Accordingly, we design SingVisio in two versions: a basic version and an advanced version. Both versions include C1 and C2 tasks. Additionally, the advanced version encompasses C3 task, catering to the needs of experts and researchers.
Fig. 2: Visual system for diffusion-based singing voice conversion. The system consists of five views. (A) Metric View shows objective evaluation results on the singing voice conversion model, allowing users to interactively explore the performance trend along diffusion steps. (B) Projection View aids users in tracking the data patterns of diffusion steps in the embedding space under different input conditions. (C) Step View provides users with the visualization of Mel spectrogram and pitch contour at one diffusion step. (D) Comparison View facilitates users to compare voice conversion results among different diffusion steps or singers. (E) Control Panel enables users to select various comparison modes and choose different source and target singers to visually understand and analyze the model behavior. The red annotations provide explanations for the patterns or components.
@@ 4.2 Analytical Tasks
Our system is a visualization system designed specifically for diffusion-based SVC tasks. Diffusion-based SVC itself involves two aspects: in the realm of machine learning, it involves the diffusion generative model, and in the field of audio signal processing, it pertains to SVC. Therefore, the analytical tasks supported by our system can be divided into two major categories. In the aspect of machine learning, particularly in the diffusion model, to investigate the evolution and quality of the generated result from each step in the diffusion generation process, our system should support the following two tasks.
@@ 5 Explainer System
Fig. 2 shows the overview of the explainer system which consists of five components: Control Panel allows users to modify mode and choose data for visual analysis; Step View provides users with an overview of the diffusion generation process; Comparison View makes it easy for users to compare converted results between different conditions; Projection View helps users observe the trajectory of diffusion steps with or without conditions; Metric View displays objective metrics evaluated on the diffusion-based SVC model, enabling users to interactively examine metric trends across diffusion steps.
@@ 5.1 Control Panel
    Step Comparison This mode primarily focuses on step-wise comparing the diffusion steps in the generation process. It (1) provides an animation of random noise gradually refined for users to have an overview of the whole denoising process in Step View, (2) enables users to adaptively select and compare the generated results from different diffusion steps in Comparison View.
  • 2.
@@
  • 3.
    Source Singer Comparison This mode focuses on the pair-wise comparison of converting two different source singers’ audio with the same song to the same target singer. It (1) allows users to select two different source singers, a source song and a target singer, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the two source audio and the target audio in the Comparison View, (3) presents two conversion animations wherein random noise undergoes gradual refinement to transform into the singing voice of the target singer in the Step View. This mode is only available in the advanced version.
  • 4.
    Song Comparison This mode focuses on the pair-wise comparison of converting two different source audios that are derived from the same singer but contain different songs to the same target singer. It (1) allows users to select a source singer, a target singer but two songs, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the two source singers’ audio and the target singer’s audio in the Comparison View, (3) supplies two conversion animations illustrating the progressive refinement of random noise into the singing voice of the target in the Step View. This mode is only available in the advanced version.
  • 5.
    Target Singer Comparison This mode focuses on the pair-wise comparison of converting the same source singing voice (also means the same song) to two different target singers. It (1) enables users to select a source singer and a source song, but two target singers, (2) provides the details (including Mel spectrogram, pitch contour, and audible audio) of the source singer’s audio, and two target singers’ audio in the Comparison View, (3) provides the two corresponding conversion animations of random noise gradually refined to the target singer singing voice in the Step View. This mode is only available in the advanced version.
@@
Projection Embedding A drop-down box to enable users to choose different projection embeddings from different layers. Then, the system displays 2D t-SNE visualization results of the high-dimensional diffusion steps in the Projection View. Specifically, the projection embedding can be the diffusion steps, the combined embeddings of the step and noise, or step, noise and conditions. These embeddings can come from the first, middle, or final residual layer in the diffusion model, as illustrated in Fig. 7 of C.
Components Two checkboxes, labeled ’F0 contour’ and ’Frequency,’ allow users to control the display of these components in the Mel spectrogram. Additionally, the frequency bar lets users adjust the frequency range for display.
Step Controller The Step Controller includes (1) a step slider to smoothly control the diffusion step, (2) a tool-tip to display or input a specific step number, and (3) a button named ‘Pin’ that enables users to add a specific step’s generated result in the Comparison View.
@@ 5.2 Step View
This view enables users to visualize the whole generation process of diffusion in the context of SVC tasks, which means users can observe how the spectral characteristics change over time as noise is subsequently removed, leading to the desired SVC. Specifically, it can be observed that the Mel spectrogram transitions from being completely noisy to gradually becoming clearer, and the fundamental frequency curve also transforms from scattered points into a smooth curve. The audio also undergoes a process of gradual optimization from being pure noise to having improved sound quality and intelligibility.
@@
Fig. 3: The left part is the Metric View with MCD metric selected. The right part is the corresponding “Metric Curve over Diffusion Step” for the best-performing sample on the MCD metric. The red annotation in the right part explains the tendencies of metric curves.
@@ 5.3 Comparison View
To facilitate a more convenient and detailed observation of the intermediate results generated by the diffusion model, we introduce a Comparison View. Moreover, the comparison view differs between the basic and advanced versions. In the basic version, the comparison view initially displays a step comparison matrix, highlighting differences in Mel spectrograms between pairs of steps in the diffusion model, as shown in Fig. 2. Darker colors in the step comparison matrix indicate larger differences, while lighter colors represent smaller ones. Users can add specific steps to the matrix using the pin feature in the control panel or by clicking data points in the projection view. By clicking on the comparison matrix, Mel spectrograms and audio of the corresponding two steps can be displayed in the comparison view for detailed comparison. In the advanced version, we directly display three Mel spectrograms from three steps by default. Besides, users can select any step to replace the displayed three steps. It enables users to compare differences among three steps, broadening the scope of comparison.
@@
It is noted that along with the Mel spectrogram, the corresponding audible audio, and fundamental frequency (F0) contour are also displayed in the Comparison View. All the information related to a clip of audio forms a basic block referred to as the “basic display unit”, as shown in the below two Mel spectrograms in the comparison view in Fig. 2 On this basic display unit, we can observe the range of the F0 and the pattern of the F0 contour. Through the brush operation, we can synchronously magnify all Mel spectrograms illustrated in this view, thus enabling a more detailed comparison and examination of the spectral differences. When there is more than one basic display unit, users can select the checkboxes in the top left corner of any two basic display units. The page will then pop up the visualization of the difference in the Mel spectrogram between these two basic display units, allowing for a clearer and more convenient comparison. Specifically, the differences are represented by colors. Warmer colors like reds and oranges signify larger differences, while cooler colors like blues and greens represent smaller differences between the two selected Mel spectrograms. This visualization aids in identifying which parts of the Mel spectrogram are significantly refined during the step-by-step generation process, highlighting areas that may require further investigation by algorithm researchers.
@@
Furthermore, the components displayed in the Comparison View differ between the Step Comparison Mode and the Condition Comparison Mode. The Step Comparison Mode is primarily used to compare the results of different diffusion steps. In this mode, the Comparison View will display basic display units from three different steps, based on the steps selected by the user. The relative position of different basic display units (corresponding to different steps) can be directly adjusted by dragging. On the other hand, the Condition Comparison Mode sports a similar layout but mainly compares the results of SVC under different conditions. In this mode, the Comparison View primarily displays the basic display units corresponding to different audios of the source and target singers selected by users. Additionally, in Metric Comparison Mode, the Comparison View illustrates the metric curve over the diffusion step, which is described in Section 5.5.
@@ 5.4 Projection View
High-dimensional hidden features can be challenging to interpret directly. t-SNE reduces the dimensionality by projecting the hidden feature embedding into a lower-dimensional space, allowing researchers to gain insights into the intricate structure and relationships within the high-dimensional space. By projecting high-dimensional step embeddings in the diffusion model into a lower-dimensional space, t-SNE reveals patterns and trajectories of the diffusion steps, enabling a visual exploration of the dynamic evolution of the diffusion process. Consequently, we design Projection View to present the two-dimensional space obtained by projecting high-dimensional diffusion step embeddings (i.e., the step features in Fig. 7 of C), as shown in Fig. 2. Each point represents a diffusion step, and all 1000 diffusion steps together form a trajectory in space. The boundary of this trajectory is highlighted with a gradient color scheme ranging from blue to red, reflecting the progression of the generation process. Users can hover their mouse over the points and slide to inspect the step trajectory. While sliding the mouse, users can simultaneously observe the SVC results transition from a coarse state to a fine state in Step View. By scrolling the mouse wheel, they can zoom in or out on the points in the space to explore the distribution of the data points. By clicking on a specific step point, a basic display unit corresponding to the step will be added into the Comparison View.
As described in Section 5.1, the drop-down menu of projection embedding in the control panel provides multiple projection embedding sources, including not only the vanilla diffusion step but also the combination of the diffusion step with noise and condition, as indicated by the red solid dots in Fig. 7 of C. By examining the projection embedding results of combining diffusion step with noise and condition, users can compare the differences in diffusion step trajectories under different condition scenarios. Additionally, we propose a novel sampling strategy called interval clustering center sampling. The specific steps are as follows: (a) Set the number of steps to be sampled, denoted as $S$. (b) Divide the total interval into T/S sub-intervals, each sampling one sample. (c) Perform k-means clustering on all samples within each sub-interval to find the central point, and then calculate the distance from all samples to this center, and finally select the sample closest to the center as the representative sample. The clustering approach ensures that each sub-interval selection considers global temporal embedding information. Furthermore, by clustering and selecting the step closest to the center, the chosen steps are highly representative.
@@ 5.5 Metric View
The Metric View represents an average of all samples within the testing data pool, providing a comprehensive overview of the model performance with five objective metrics. Upon hovering over a particular metric, the system automatically identifies and selects the best-performing sample. This selection triggers a detailed visualization of the diffusion step for that sample within the Projection View. Additionally, for the chosen sample, the system dynamically computes and displays evaluation metrics, which are then used to plot the “Metric Curve over Diffusion Step” in Comparison View, as shown in Fig. 3.